OpenStudy (anonymous):

how do i multiply 17/19 by 26

OpenStudy (anonymous):

\[\frac{17\times 26}{19}\]

OpenStudy (anonymous):

^ And them simplify if possible.

OpenStudy (anonymous):

unlikely since 17 and 19 are prime

OpenStudy (anonymous):

That's why I said "if possible", so she'd know for future problems. :)

OpenStudy (anonymous):

I'm confused on what the answer is still

OpenStudy (anonymous):

What's 17 * 26

OpenStudy (anonymous):

442

OpenStudy (anonymous):

So then you get 442/19 right?

OpenStudy (anonymous):

yea i got a decimal but i can't use a decimal in my answer

OpenStudy (anonymous):

No, don't divide it, turn it into an improper fraction

OpenStudy (anonymous):

I dont know how to do that

OpenStudy (anonymous):

Well, yes, divide it, but turn that decimal into a fraction

OpenStudy (anonymous):

17 * 26 = 442 442/19 = 23.26315789473684 Round that decimal to ----- 23.3 Your fraction would be 23 3/10

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You're welcome :)